Question 19

A normally distributed population with 1000 observations has a mean of 100 and a standard deviation 10. A sample of size 200 is randomly drawn from the population.

What is the standard error of the sampling distribution?
(1) 10
(2) 0.707
(3) 0.05
(4) 0.1
(5) 0.316
1
Expert's answer
2020-03-03T11:00:29-0500

"SE=\frac{\sigma}{\sqrt{n}}"

"n=200, \sigma=10"

"SE=\frac{10}{\sqrt{200}}= 0.707"